10 years from now the sum of the ages of A and B is equal to 50. 6 years ago, the difference of their ages is equal to 6.how old A and B?
Answer by ewatrrr(24785)

A + B + 20 = 50
A - B - 12 = 6
....
A + B = 30
A - B = 18
2A = 48
A = 24 and B = 6
And...checking
10 years from now...
34 + 16 = 50